OpenStudy (anonymous):

Find a ⋅ b. a = <7, 4>, b = <3, 5>

Parth (parthkohli):

Dot product?

OpenStudy (anonymous):

@ParthKohli yes

Parth (parthkohli):

\[a \cdot b\]\[= \langle 7,4\rangle \cdot \langle 3,5\rangle \]\[= 7\times 3 + 4\times 5\]\[= 21 + 20\]\[= 41\]

OpenStudy (anonymous):

@ParthKohli that's all I do?

Parth (parthkohli):

Yes, that's all.

OpenStudy (anonymous):

@ParthKohli thanks!

Parth (parthkohli):

No problem!